# Slow-Fast Policy Optimization: Reposition-Before-Update for LLM Reasoning

### v1

@@ 1 Introduction
Despite these successes, GRPO inherits structural inefficiencies that make training fragile in LLM reasoning. During the early stages of training, when rollouts are weak or uninformative, stochastic rewards induce high-variance gradients that destabilize updates. Prior studies corroborate this issue: early rollouts often fail to provide useful signals (Zheng et al., 2025), reward shaping can trigger sudden collapse (Dai et al., 2025), and noisy group normalization further impairs learning under imbalanced responses (Shen et al., 2025). Although group normalization partially mitigates variance, GRPO still applies only a single gradient update per batch, discarding information that multiple inner steps could exploit (Schulman et al., 2017, Engstrom et al., 2020). As a result, reasoning RL training tends to be sample-inefficient, with early exploration fragile and progress highly sensitive to noise and rollout quality. Although the original GRPO paper (Shao et al., 2024) introduces Iterative GRPO to reuse rollout data, this is primarily designed for updating the reward model. We find that applying this off-policy update to the policy model can degrade performance, and how to use it effectively to build strong reasoning abilities in LLMs remains an open question.
@@ 4.2.1 Math Reasoning Benchmarks
Figure 2: Average validation accuracy of different base models throughout the learning process.
@@ 5 Analysis and Ablation Study
Impact of $\alpha$ and $K$. As discussed in Sec. 3, the hyperparameters $\alpha$ and $K$ jointly determine the stability of SFPO. To better understand their impacts, we study two settings: small K=3K{=}3 and large K=7K{=}7, under varying $\alpha$, as shown in Fig. 5. When $K$ is small, SFPO remains stable across different $\alpha$ values and consistently outperforms GRPO. This aligns with our theoretical intuition that performing multiple inner updates on the same rollout batch, followed by reposition, effectively reduces gradient noise and produces a more reliable update direction. However, when $K$ is large, the fast weights drift substantially from the original parameters. A large $\alpha$ then amplifies this mismatch by pulling the slow weights too aggressively toward unstable fast trajectories, injecting noise and causing performance drops. As shown in Fig. 5(b), a smaller $\alpha$ mitigates this drift and restores stability—confirming the interaction between $K$ and $\alpha$.
Figure 5: Average training accuracy of different base models throughout the learning process.
@@
Interpolation Against Off-Policy Overfitting. When $\alpha=1$ (no interpolation), performance initially rises but steadily declines as training progresses. Without interpolation, the model rapidly adapts by overfitting to small batches of rollouts—yielding short-term gains but injecting growing noise into gradient updates, which causes instability and long-term degradation. This effect is amplified when $K$ is large, as shown in Fig. 5(b), where the fast weights drift substantially and the slow weights fully adopt these noisy trajectories, leading to sharp performance drops. In contrast, when $\alpha$ is small, interpolation effectively anchors the update near the on-policy region, mitigating distributional drift and stabilizing training. However, overly small $\alpha$ may under-utilize the stabilized fast trajectory, slightly slowing early-stage improvement. Overall, moderate $\alpha$ achieves the best balance—leveraging variance reduction from Stage I while maintaining stability through controlled repositioning.
@@
Necessity of Entropy Control. We evaluate the effect of adaptive entropy control (EC) on SFPO using the DS-Qwen-7B model. As shown in Fig. 7, removing EC causes a noticeable accuracy drop after roughly 100 steps. This degradation coincides with a rapid divergence of entropy loss, suggesting that the policy becomes unstable and overfits to noisy rollouts. These results highlight entropy control as a key factor for maintaining the stability and reliability of SFPO. For simplicity, we implement entropy control through a lightweight scheduling of $\alpha$. As shown in Alg. 1, once a predefined entropy-trigger is activated, $\alpha$ is set to 0, reverting SFPO to standard GRPO for subsequent steps. This simple mechanism effectively stabilizes training while retaining the efficiency gains of our fast–reposition–slow framework.
@@ 6 Related Works
Despite focusing on designing novel training pipelines, a complementary line of work (Ivison et al., 2025, Xia et al., 2024, Muennighoff et al., 2025, Ye et al., 2025) improves the efficiency of LLM training through data filtering. One direction focuses on pruning data for supervised fine-tuning (Xia et al., 2024, Chen et al., 2023, Ivison et al., 2022). Another direction targets reinforcement learning, where studies (Muldrew et al., 2024, Liu et al., 2024, Das et al., 2024, Li et al., 2025, Fatemi et al., 2025, Wang et al., 2025) show that GRPO requires only a small subset of the training data to improve reasoning ability. However, these methods largely optimize which data is used rather than how updates are performed. SFPO is orthogonal: it assumes no change to the data pipeline but restructures the policy update itself. By converting one-shot updates into a fast–reposition–slow trajectory, SFPO reduces variance and stabilizes learning, thereby yielding higher data efficiency even under random rollout sampling.
Iterative GRPO.
Iterative GRPO was introduced in (Shao et al., 2024) to reuse rollout data, with a primary focus on reward modeling. Mroueh (2025) provided a theoretical analysis under verifiable rewards, reformulating GRPO as a KL-regularized contrastive loss and proving its iterative dynamics amplify success rates relative to the reference policy. Mroueh et al. (2025) examined both on- and off-policy variants, deriving reward improvement conditions, introducing clipped surrogate objectives to stabilize training, and empirically showing off-policy GRPO can match or exceed on-policy performance. Unlike these modifications, SFPO preserves the on-policy GRPO update but restructures one-step optimization into a fast-reposition-slow framework to tackle sample inefficiency and stabilize updates.

### v2

@@ 1 Introduction
Despite these successes, GRPO inherits structural inefficiencies that make training fragile in LLM reasoning. During the early stages of training, when rollouts are particularly weak or uninformative, stochastic rewards induce high-variance gradients that destabilize updates (Yu et al., 2025a, Zheng et al., 2025, Dai et al., 2025, Shen et al., 2025). Compounding this, common practice in GRPO implementations is to use a single optimization step per rollout batch, and such a one-shot update tends to produce a noisier, less reliable step update direction, especially at the early training stages, thereby underutilizing each batch and discarding information that could be further exploited (Schulman et al., 2017, Engstrom et al., 2020). Although the original GRPO paper (Shao et al., 2024) explored reusing rollout data, our findings suggest that simply applying this off-policy update to the policy model can degrade performance as training proceeds, and how to use it effectively to build strong reasoning abilities in LLMs with satisfying sample efficiency remains an open question.
@@ 4.2.1 Math Reasoning Benchmarks
Figure 2: Average validation accuracy of different base models throughout the training process.
@@ 5 Analysis and Ablation Study
Impact of $\alpha$ and $K$. As discussed in Sec. 3, the hyperparameters $\alpha$ and $K$ jointly determine the stability of SFPO. To better understand their impacts, we study two settings: small K=3K{=}3 and large K=7K{=}7, under varying $\alpha$, as shown in Fig. 5. When $K$ is small, SFPO remains stable across different $\alpha$ values and consistently outperforms GRPO. This aligns with our theoretical intuition that performing multiple inner updates on the same rollout batch, followed by reposition, effectively reduces gradient noise and produces a more reliable update direction. However, when $K$ is large, the fast weights drift substantially from the original parameters. A large $\alpha$ then amplifies this mismatch by pulling the slow weights too aggressively toward unstable fast trajectories, injecting noise and causing performance drops. As shown in Fig. 5(b), a smaller $\alpha$ mitigates this drift and restores stability, confirming the interaction between $K$ and $\alpha$.
Figure 5: Average training accuracy of different settings throughout the training process. (a): Small k=3 with varying values of $\alpha$. (b): Large k=7 with varying values of $\alpha$. (c): Varying values of k with fixed $\alpha=0.8$. (d): Impact of the existence of stage III.
@@
Interpolation Against Off-Policy Overfitting. When $\alpha=1$ (no interpolation), performance initially rises but steadily declines as training progresses. Without interpolation, the model rapidly adapts by overfitting to small batches of rollouts, yielding short-term gains but injecting growing noise into gradient updates, which causes instability and long-term degradation. This effect is amplified when $K$ is large, as shown in Fig. 5(b), where the fast weights drift substantially and the slow weights fully adopt these noisy trajectories, leading to sharp performance drops. In contrast, for small $\alpha$, interpolation keeps the update closer to the on-policy region, mitigating distributional drift and stabilizing training. However, if $\alpha$ is overly small and $K$ is small, the algorithm under-utilizes the stabilized fast trajectory, slightly slowing early improvements. While for large $K$, a smaller $\alpha$ is preferable, where interpolation effectively counteracts the severe off-policy drift induced by many inner steps. Overall, interpolation serves as a simple, effective regularizer against the potential off-policy overfitting introduced by Stage I.
@@
Necessity of Entropy Control. We evaluate the effect of adaptive entropy control (EC) on SFPO using the DS-Qwen-7B model. As shown in Fig. 7, removing EC causes a noticeable accuracy drop after roughly 100 steps. This degradation coincides with a rapid divergence of entropy loss, suggesting that the policy becomes unstable and overfits to noisy rollouts. These results highlight entropy control as a key factor for maintaining the stability and reliability of SFPO. For simplicity, we implement entropy control through a lightweight scheduling of $\alpha$. As shown in Alg. 1, once a predefined entropy-trigger is activated, $\alpha$ is set to 0, reverting SFPO to standard GRPO for subsequent steps. This simple mechanism effectively stabilizes training while retaining the efficiency gains.
@@ 6 Related Works
Despite focusing on designing novel training pipelines, a complementary line of work (Ivison et al., 2025, Xia et al., 2024, Muennighoff et al., 2025, Ye et al., 2025) improves the efficiency of LLM training through data filtering. One direction focuses on pruning data for supervised fine-tuning (Xia et al., 2024, Chen et al., 2023, Ivison et al., 2022). Another direction targets reinforcement learning, where studies (Muldrew et al., 2024, Liu et al., 2024, Das et al., 2024, Li et al., 2025, Fatemi et al., 2025, Wang et al., 2025) show that GRPO requires only a small subset of the training data to improve reasoning ability. However, these methods largely optimize which data is used rather than how updates are performed. SFPO is orthogonal: it assumes no change to the data pipeline but restructures the policy update itself. By converting one-shot updates into a fast–reposition–slow trajectory, SFPO reduces variance and stabilizes learning.
Off-policy GRPO.
Original iterative GRPO  (Shao et al., 2024) tries to update policy model for multiple times with the same batch of rollout data. Mroueh (2025) provided a theoretical analysis for this off-policy policy gradient operation under verifiable rewards, reformulating GRPO as a KL-regularized contrastive loss and proving its iterative dynamics amplify success rates relative to the reference policy. Mroueh et al. (2025) examined both on- and off-policy variants, deriving reward improvement conditions, introducing clipped surrogate objectives to stabilize training, and empirically showing off-policy GRPO can match or exceed on-policy performance. Unlike these modifications, SFPO preserves the on-policy GRPO update but restructures one-step optimization into a fast-reposition-slow framework to tackle sample inefficiency and stabilize updates.